solve by the substitution method
5m+n=7
m-6n=51
The solution is?

Answer by algebrahouse.com(1659)

solve by the substitution method
5m + n = 7-------> n = -5m + 7 {subtracted 5m from both sides}
m - 6n = 51

m - 6n = 51 {bottom original equation}
m - 6(-5m + 7) = 51 {substituted -5m + 7, in for n, into bottom equation}
m + 30m - 42 = 51 {used distributive property}
31m - 42 = 51 {combined like terms}
31m = 93 {added 42 to both sides}
m = 3 {divided both sides by 31}

n = -5m + 7
n = -5(3) + 7 {substituted 3, in for m, into n = -5m + 7}
n = -15 + 7 {multiplied -5 by 3}
n = -8 {combined like terms}

m = 3 and n = -8
www.algebrahouse.com